OpenStudy (anonymous):

tan(theta/2)−sin(theta)= 0 solve for all possible soultions

OpenStudy (campbell_st):

well the obvious solution is when then angle theta = 0 and then every 2pi after that the equation will be tan(0/2) - sin(0) = 0 the other solution is when tan (x/2) = sin(x) this occurs when x = pi/2 tan(pi/2/2) = tan(pi/4) = 1 and sin(pi/2) = 1 and the 3rd solution when tan(x/2) = -1 and sin(x) = -1 hope this helps

OpenStudy (campbell_st):

so 3 of the solutions are 0, pi/2 and 3pi/2 then they all repeat every 2pi